2/5(x+1)=8/15; (1/6, 1/3, 1/2, 2/3). () are brakets
Answer by edjones(8007)

2/5(x+1)=8/15
x+1=4/3 multiply each side by 5/2
x=4/3 - 1
x=1/3
.
Ed